# Synthetic Tabular Data Validation: A Divergence-Based Approach

### v1

@@ Abstract
The ever-increasing use of generative models in various fields where tabular data is used highlights the need for robust and standardized validation metrics to assess the similarity between real and synthetic data. Current methods lack a unified framework and rely on diverse and often inconclusive statistical measures. Divergences, which quantify discrepancies between data distributions, offer a promising avenue for validation. However, traditional approaches calculate divergences independently for each feature due to the complexity of joint distribution modeling. This paper addresses this challenge by proposing a novel approach that uses divergence estimation to overcome the limitations of marginal comparisons. Our core contribution lies in applying a divergence estimator to build a validation metric considering the joint distribution of real and synthetic data. We leverage a probabilistic classifier to approximate the density ratio between datasets, allowing the capture of complex relationships. We specifically calculate two divergences: the well-known Kullback-Leibler (KL) divergence and the Jensen-Shannon (JS) divergence. KL divergence offers an established use in the field, while JS divergence is symmetric and bounded, providing a reliable metric. The efficacy of this approach is demonstrated through a series of experiments with varying distribution complexities. The initial phase involves comparing estimated divergences with analytical solutions for simple distributions, setting a benchmark for accuracy. Finally, we validate our method on a real-world dataset and its corresponding synthetic counterpart, showcasing its effectiveness in practical applications. This research offers a significant contribution with applicability beyond tabular data and the potential to improve synthetic data validation in various fields.
@@ 1 Introduction
Following the impressive advances in GMs for data generation, a critical aspect emerges: ensuring the quality and effectiveness of synthetic data. Replicating statistical properties, accuracy, consistency, and domain-specific relevance of the source data is essential. This challenge becomes especially pronounced as GMs gain traction within engineering design communities. The need for robust and standardized validation metrics becomes paramount. The current literature highlights a significant gap in this area. Although there are standardized evaluation metrics for the generation of synthetic images [4], [5], and the generation of text [6], [7], measuring the quality of synthetic tabular data presents unique challenges. Unlike image data, qualitative evaluation through visual inspection is not feasible. Additionally, relying solely on expert insight can be highly inefficient. The existing landscape for tabular data validation often focuses primarily on the efficacy or utility in machine learning tasks [8][9]. However, a consistent approach to assess the similarity between synthetic and real tabular data remains sparse. Studies employ a diverse set of metrics, including pairwise correlation difference[10], support coverage[11], likelihood fitness[12], alongside various statistic values[13],[14]. Reviews such as [15] attempt to categorize these techniques, but a truly standardized approach or a single metric that captures the full spectrum of statistical information from a distribution remains a challenge.
@@
This paper aims to address this gap by proposing a divergence estimator. Divergences offer a comprehensive approach to the validation of similarity. These measures quantify the discrepancy between two probability distributions, effectively capturing the differences between real and synthetic data. Unlike traditional methods that often focus on individual attributes [16][17], divergences can consider the joint distribution of all attributes, providing a more holistic view of the data. However, it is important to acknowledge the challenge associated with joint distributions. Modeling the intricate relationships between all features in a dataset can be computationally expensive and intractable. Although some existing papers explore divergences or distances for validation, they often apply them to single attributes because of this limitation [11][18][19]. Crucially, even if individual features appear similar, the joint distribution, which captures the relationships between them, may differ significantly [20]. This underscores the importance of considering the joint distribution for accurate assessment. Our proposed divergence estimator addresses this challenge by offering a way to efficiently capture these complex joint distributions. This allows us to compute divergences between real and synthetic data, enabling a more comprehensive and robust evaluation of synthetic data quality. The advantages of using divergences for validation remain the same:
@@
  • Sensitivity to Complex Distributions: Divergences can capture the full complexity of real-world data distributions, including nonlinear relationships and rare events, which traditional methods might miss.
@@
  • Interpretability: Divergences can be easily interpreted and visualized, providing valuable insights into the discrepancies between real and synthetic data.
@@
  • Divergence-based Similarity Estimation: Our approach leverages divergence estimation to quantify the discrepancy between the joint probability distributions of real and synthetic tabular data. This approach goes beyond traditional validation metrics, which often focus on marginal statistics.
@@
  • KL Divergence and JS Divergence for Validation: We acknowledge the widespread use of Kullback-Leibler (KL) divergence in information theory (often referred to as relative entropy) for divergence calculations. However, we recognize its limitations: KL divergence is not a true distance in terms of mathematics [22] due to its asymmetry and the potential for infinite values in certain scenarios. To address this, we propose the inclusion of the Jensen-Shannon (JS) divergence, which offers desirable properties of being bounded and symmetric, thus defining a proper distance metric. This bounded metric is crucial to assess the quality of synthetic data generation.
@@
  • Broad Applicability: Due to the adaptability of our divergence-based validation method, it can be applied to various types of data beyond tabular formats. This has the potential to significantly impact numerous fields such as healthcare, finance, and manufacturing. The adoption of our approach can lead to substantial improvements in the accuracy and reliability of synthetic data validation in various domains.
@@ 2.1 Divergence definition
Let us consider two probability distributions, $p(x)$ and $q(x)$, for a random variable $x$. In the context of probability distributions, divergence quantifies the dissimilarity between $p(x)$ and $q(x)$. In essence, it measures how different these distributions are in terms of the probabilities they assign to each possible value of $x$. Divergences play a crucial role in various fields, including Machine Learning (ML), as they provide a way to compare and analyze the behavior of different probability distributions.
@@
$D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}=\frac{1}{2}D_{\mathbb{KL}}\big{(}p(x)\|m(x)\big{)}+\frac{1}{2}D_{\mathbb{KL}}\big{(}q(x)\|m(x)\big{)}.$ (3)
The common reference distribution, $m(x)$, is typically chosen as the midpoint between the two original distributions, such as the average: $m(x)=\big{(}p(x)+q(x)\big{)}/2$. Importantly, the JS divergence is inherently bounded. When employing a base-$2$ logarithm, the maximum divergence value is $1$.
@@
$\approx\frac{1}{2L}\sum_{i=1}^{L}\log\frac{p\left(x_{i}\right)}{m\left(x_{i}\right)}+\frac{1}{2L}\sum_{i=1}^{L}\log\frac{q\left(\tilde{x}_{i}\right)}{m\left(\tilde{x}_{i}\right)}=\frac{1}{2L}\sum_{i=1}^{L}\log\left(\frac{2p\left(x_{i}\right)}{p\left(x_{i}\right)+q\left(x_{i}\right)}\right)+\frac{1}{2L}\sum_{i=1}^{L}\log\left(\frac{2q\left(\tilde{x}_{i}\right)}{p\left(\tilde{x}_{i}\right)+q\left(\tilde{x}_{i}\right)}\right),$
@@ 2.2 Estimating density ratio using probabilistic classification
Density ratio estimation focuses on estimating the ratio $r^{*}(x)$ of two probability densities, $p(x)$ and $q(x)$, based solely on samples drawn from these distributions. The ratio is therefore defined as
@@
Estimating (1) and (3) is based on determining the unknown density ratio $r^{*}$. This section establishes a connection between the density ratio of two distributions $p(x)$ and $q(x)$, and a probabilistic classifier that optimally distinguishes samples drawn from these distributions. Consider a scenario in which we possess samples from both distributions, $p(x)$ and $q(x)$, with each sample labeled according to its origin. Using a classifier to estimate the class-membership probabilities for each sample, we can derive an estimator for the density ratio. Let $X_{p}={x^{(1)},x^{(2)},...,x^{(M_{p})}}$ and $X_{q}={\tilde{x}^{(1)},\tilde{x}^{(2)},...,\tilde{x}^{(M_{q})}}$ represent sets of real and synthetic data samples, respectively, where $M_{p}$ and $M_{q}$ denote the corresponding sample sizes.
@@
$r^{*}(x)=\frac{p(x)}{q(x)}=\frac{\mathcal{P}(x\mid y=1)}{\mathcal{P}(x\mid y=0)}=\left(\frac{\mathcal{P}(y=1\mid x)\mathcal{P}(x)}{\mathcal{P}(y=1)}\right)\left(\frac{\mathcal{P}(y=0\mid x)\mathcal{P}(x)}{\mathcal{P}(y=0)}\right)^{-1}=\frac{\mathcal{P}(y=0)}{\mathcal{P}(y=1)}\frac{\mathcal{P}(y=1\mid x)}{\mathcal{P}(y=0\mid x)}.$ (7)
Assuming equal prior probabilities for the source distributions (i.e., $\mathcal{P}(y=0)=\mathcal{P}(y=1)$), the ratio of marginal probabilities cancels out, leaving:
@@
$\sigma^{-1}(z)=\ln\left(\dfrac{z}{1-z}\right)$ (9)
where $z$ represents any probability. The logit function has the property that the log odds of two probabilities equal the log of their ratio. In our case, applying the logit function to both the numerator and denominator of the expression for 8 results in the log ratio of the original probabilities:
$\ln\left(\dfrac{\mathcal{P}(y=1\mid x)}{\mathcal{P}(y=0\mid x)}\right)=\sigma^{-1}\left(\mathcal{P}(y=1\mid x)\right)-\sigma^{-1}\left(\mathcal{P}(y=0\mid x)\right)$ (10)
@@ 3 Methodology
This research uses a neural network (NN) classifier, denoted $D_{\theta}$ with parameters $\theta$, to approximate the posterior probability $P(y=1\mid x)$. This network acts as a discriminator to classify input samples originating from the $p(x)$ or $q(x)$ distributions. The network architecture is designed to capture the discriminative features that distinguish these distributions effectively.
@@
$r_{\theta}(x)=\exp\big{[}\sigma^{-1}\left(D_{\theta}(x)\right)\big{]}\approx\exp\big{[}\sigma^{-1}(\mathcal{P}(y=1\mid x))\big{]}=r^{*}(x).$ (13)
@@
$\mathcal{L}(\theta)=-\mathbb{E}_{p(x)}\big{[}\log\mathcal{D}_{\theta}(x)\big{]}-\mathbb{E}_{q(x)}\big{[}\log(1-\mathcal{D}_{\theta}(x))\big{]}.$ (14)
Once trained, the class probability estimator can be used to construct MC estimates of the KL and JS divergences. The KL divergence between $p(x)$ and $q(x)$ can be estimated as
$D_{\mathbb{KL}}\big{(}p(x)\|q(x)\big{)}$ (15)
$=\mathbb{E}_{p(x)}\big{[}\log r^{*}(x)\big{]}\approx\frac{1}{L}\sum_{i=1}^{L}\log r^{*}(x_{i})\approx\frac{1}{L}\sum_{i=1}^{L}\log r_{\theta}(x_{i})=\frac{1}{L}\sum_{i=1}^{L}\sigma^{-1}\big{(}\mathcal{D}_{\theta}\left(x_{i}\right)\big{)},$
@@
$\approx\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}2\mathcal{D}_{\theta}(x_{i})\big{)}+\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}2-2\mathcal{D}_{\theta}(\tilde{x}_{i})\big{)}=\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}\mathcal{D}_{\theta}(x_{i})\big{)}+\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}1-\mathcal{D}_{\theta}(\tilde{x}_{i})\big{)},$
@@
$2\cdot D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}-\log 4\geq\sup_{\theta}\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}+\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}.$ (18)
Following this analysis, we can arrive at the desired inequality. However, for a more convenient form, we can manipulate the inequality by negating both sides.
$-2\cdot D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}+\log 4\leq-\sup_{\theta}\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}+\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}$ (19)
$=\inf_{\theta}-\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}-\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}=\inf_{\theta}\mathcal{L}(\theta).$
@@
Figure 2: Generative model for divergence estimation between real and synthetic data. The generative model learns an approximation of $N$ samples from real data $x_{r}$, denoted $x_{s}$. Subsequently, $M$ samples are drawn from each distribution to train the divergence estimation discriminator. Finally, $L$ samples from each distribution are used to estimate the divergence between real and synthetic data.
This study evaluates the dissimilarity between real and synthetic data generated by a Generative Model (GM). In this context, minimizing the divergence between the real and synthetic data distributions is crucial. Ideally, the divergence should approach zero, indicating that the generated data become virtually indistinguishable from the real data. This interchangeability allows the use of synthetic data in various applications where real data might be scarce or sensitive. From the discriminator’s perspective, achieving minimal divergence implies that it cannot reliably differentiate between real and synthetic samples, indicating successful data generation. For GMs, the distributions of interest become:
@@
Training data consist of $N$ samples drawn from the real data distribution, denoted $x_{r}\sim p_{R}(x)$. These data are used to train the GM, as illustrated in Fig. 2. The GM learns to approximate the real data distribution as $p_{S}(x)$, allowing us to sample from this distribution and obtain synthetic data points, $x_{s}\sim p_{S}(x)$. Following the architecture described in Fig. 1, these synthetic samples are then used along with the real data samples to train the discriminator and estimate the divergence between the real and synthetic data distributions.

### v2

@@ Abstract
The ever-increasing use of generative models in various fields where tabular data is used highlights the need for robust and standardized validation metrics to assess the similarity between real and synthetic data. Current methods lack a unified framework and rely on diverse and often inconclusive statistical measures. Divergences, which quantify discrepancies between data distributions, offer a promising avenue for validation. However, traditional approaches calculate divergences independently for each feature due to the complexity of joint distribution modeling. This paper addresses this challenge by proposing a novel approach that uses divergence estimation to overcome the limitations of marginal comparisons. Our core contribution lies in applying a divergence estimator to build a validation metric considering the joint distribution of real and synthetic data. We leverage a probabilistic classifier to approximate the density ratio between datasets, allowing the capture of complex relations. We specifically calculate two divergences: the well-known Kullback-Leibler (KL) divergence and the Jensen-Shannon (JS) divergence. KL divergence offers an established use in the field, while JS divergence is symmetric and bounded, providing a reliable metric. The efficacy of this approach is demonstrated through a series of experiments with varying distribution complexities. The initial phase involves comparing estimated divergences with analytical solutions for simple distributions, setting a benchmark for accuracy. Finally, we validate our method on a real-world dataset and its corresponding synthetic counterpart, showcasing its effectiveness in practical applications. This research offers a significant contribution with applicability beyond tabular data and the potential to improve synthetic data validation in various fields.
@@ 1 Introduction
Following the impressive advances in GMs for data generation, a critical aspect emerges: ensuring the quality and effectiveness of synthetic data. Replicating the source data’s statistical properties, accuracy, consistency, and domain-specific relevance is essential. This challenge becomes significantly pronounced as GMs gain traction within engineering design communities. The need for robust and standardized validation metrics becomes paramount. The current literature highlights a significant gap in this area. Although there are standardized evaluation metrics for the generation of synthetic images [4], [5], and the generation of text [6], [7], measuring the quality of synthetic tabular data presents unique challenges. Unlike image data, qualitative evaluation through visual inspection is not feasible. Additionally, relying solely on expert insight can be highly inefficient. The existing landscape for tabular data validation often focuses primarily on the efficacy or utility in machine learning tasks [8], [9]. However, a consistent approach to assessing the similarity between synthetic and real tabular data remains sparse. Studies employ a diverse set of metrics, including pairwise correlation difference[10], support coverage[11], likelihood fitness[12], alongside various statistic values[13],[14]. Reviews such as [15] attempt to categorize these techniques, but a genuinely standardized approach or a single metric that captures the full spectrum of statistical information from a distribution remains a challenge.
@@
This paper aims to address this gap by proposing a divergence estimator. Divergences offer a comprehensive approach to the validation of similarity. These measures quantify the discrepancy between two probability distributions, effectively capturing the differences between real and synthetic data. Unlike traditional methods that often focus on individual attributes [16], [17], divergences can consider the joint distribution of all attributes, providing a more holistic view of the data. However, it is essential to acknowledge the challenge associated with joint distributions. Modeling the intricate relations between all features in a dataset can be computationally expensive and intractable. Although some existing papers explore divergences or distances for validation, they often apply them to single attributes because of this limitation [11], [18], [19]. Crucially, even if individual features appear similar, the joint distribution, which captures their relations, may differ significantly. Estimating divergences between marginal distributions neglects the correlation between features. For example, in survival analysis databases, the risk associated with a disease is often closely related to the patient’s age [20]. Although these two variables might be generated to mimic the original distributions, this method might fail to preserve their intrinsic, possibly nonlinear, correlation. In practice, disrupting original correlations can result in synthetic data that, while statistically similar in univariate terms, diverges significantly in multivariate contexts. This discrepancy is critical because it can affect the utility and reliability of synthetic data in applications where feature interdependencies play a crucial role, such as predictive modeling in healthcare. To address this issue, correlation matrices are sometimes employed to assess relations between features [21]. However, these matrices primarily reveal linear relations, provide complex representations usually interpreted graphically, and are challenging to analyze quantitatively. This underscores the importance of considering the joint distribution for an accurate assessment. Our proposed divergence estimator addresses this challenge by efficiently capturing these complex joint distributions. This approach enables the computation of divergences between real and synthetic data, allowing for a more comprehensive and robust evaluation of the quality of synthetic data. The advantages of using divergences for validation remain the same:
@@
  • Sensitivity to Complex Distributions: Divergences can capture the full complexity of real-world data distributions, including nonlinear relations and rare events, which traditional methods might miss.
@@
  • Interpretability: Divergences can be easily interpreted and visualized, providing valuable insights into the discrepancies between real and synthetic data. They condense statistical similarity between distributions into a single scalar, facilitating quantitative interpretation of the findings.
@@
  • Divergence-based Similarity Estimation: Our approach leverages divergence estimation to quantify the discrepancy between the joint probability distributions of real and synthetic tabular data. This approach goes beyond traditional validation metrics, often focusing on marginal statistics.
@@
  • KL Divergence and JS Divergence for Validation: We acknowledge the widespread use of Kullback-Leibler (KL) divergence in information theory (often referred to as relative entropy) for divergence calculations. However, we recognize its limitations: KL divergence is not an actual distance in terms of mathematics [23] due to its asymmetry and the potential for infinite values in specific scenarios. To address this, we propose the inclusion of the Jensen-Shannon (JS) divergence, which offers the desirable properties of being bounded and symmetric, thus defining a proper distance metric. This bounded metric is crucial to assess the quality of synthetic data generation.
@@
  • Broad Applicability: Due to the adaptability of our divergence-based validation method, it can be applied to various types of data beyond tabular formats. This has the potential to significantly impact numerous fields, such as healthcare, finance, and manufacturing. Adopting our approach can substantially improve the accuracy and reliability of synthetic data validation in various domains.
In summary, our contribution diverges from the traditional emphasis on divergence estimation, which is already a well-established and robust field of study. Instead, we focus on applying divergence estimation to validate synthetic data generation. As highlighted in our literature review, a notable opportunity exists to establish a standardized validation metric in this area. We propose that estimating divergences between real and synthetic data can serve as a definitive measure of the accuracy of the data generation process. By leveraging divergence metrics, we offer a quantitative measure that is both interpretable and grounded in solid theoretical principles. This approach introduces a robust metric that enhances the validation process for synthetic data generation, ensuring that synthetic data not only resemble the original in terms of individual feature distributions but also preserve the intricate interdependencies between features. These interdependencies are crucial for the data’s authenticity and applicability in real-world scenarios. This advancement represents a significant step forward in standardizing validation practices and improving the reliability of synthetic datasets across various fields.
The structure of the paper is carefully designed to facilitate a comprehensive understanding of our proposal. It begins with Section I, an introductory section that sets the stage by highlighting the importance of robust validation metrics in the use of synthetic data across various domains. This is followed by Section II, a detailed background section that delves into the theoretical and analytical aspects of divergence measures, specifically the KL and JS divergences. Furthermore, this section discusses the density ratio estimation based on probabilistic classification, as outlined in reference [22], which is crucial for the subsequent divergence calculation. The methodology section (Section III) then defines our approach using generative models to obtain this density ratio estimation, detailing the process analytically, visually, and programmatically. The fourth section describes the experimental setup and discusses the validation results from four different experiments, each increasing in complexity, to demonstrate the efficacy of our approach. The paper concludes in Section V with a summary of our findings, highlighting key conclusions and potential future research lines. An appendix is also included, which presents additional results from our experiments, providing deeper insights into our validation process.
@@ 2.1 Divergence definition
Let us consider two probability distributions, $p(x)$ and $q(x)$, for a random variable $x$. In the context of probability distributions, divergence quantifies the dissimilarity between $p(x)$ and $q(x)$. In essence, it measures how different these distributions are regarding the probabilities they assign to each possible value of $x$. Divergences play a crucial role in various fields, including Machine Learning (ML), as they provide a way to compare and analyze the behavior of different probability distributions.
@@
$D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}$ (3)
$=\frac{1}{2}D_{\mathbb{KL}}\big{(}p(x)\|m(x)\big{)}+\frac{1}{2}D_{\mathbb{KL}}\big{(}q(x)\|m(x)\big{)}.$
The common reference distribution, $m(x)$, is typically chosen as the midpoint between the two original distributions, such as the average: $m(x)=\big{(}p(x)+q(x)\big{)}/2$. Notably, the JS divergence is inherently bounded. When employing a base-$2$ logarithm, the maximum divergence value is $1$.
@@
$\approx$ $\frac{1}{2L}\sum_{i=1}^{L}\log\frac{p\left(x_{i}\right)}{m\left(x_{i}\right)}+\frac{1}{2L}\sum_{i=1}^{L}\log\frac{q\left(\tilde{x}_{i}\right)}{m\left(\tilde{x}_{i}\right)}$
$=$ $\frac{1}{2L}\sum_{i=1}^{L}\log\left(\frac{2p\left(x_{i}\right)}{p\left(x_{i}\right)+q\left(x_{i}\right)}\right)$
$+\frac{1}{2L}\sum_{i=1}^{L}\log\left(\frac{2q\left(\tilde{x}_{i}\right)}{p\left(\tilde{x}_{i}\right)+q\left(\tilde{x}_{i}\right)}\right),$
@@ 2.2 Estimating density ratio using probabilistic classification
Density ratio estimation focuses on estimating the ratio $r^{*}(x)$ of two probability densities, $p(x)$ and $q(x)$, based solely on samples drawn from these distributions. This field has seen significant theoretical development, particularly regarding complex ratio estimators and their convergence properties [24]. The ratio is therefore defined as
@@
Estimating (1) and (3) is based on determining the unknown density ratio $r^{*}$. This section establishes a connection between the density ratio of two distributions, $p(x)$ and $q(x)$, and a probabilistic classifier that optimally distinguishes samples drawn from these distributions. Consider a scenario in which we possess samples from both distributions, $p(x)$ and $q(x)$, with each sample labeled according to its origin. Using a classifier to estimate the class-membership probabilities for each sample, we can derive an estimator for the density ratio. Let $X_{p}={x^{(1)},x^{(2)},...,x^{(M_{p})}}$ and $X_{q}={\tilde{x}^{(1)},\tilde{x}^{(2)},...,\tilde{x}^{(M_{q})}}$ represent sets of real and synthetic data samples, respectively, where $M_{p}$ and $M_{q}$ denote the corresponding sample sizes.
@@
$r^{*}(x)$ $=\frac{p(x)}{q(x)}=\frac{\mathcal{P}(x\mid y=1)}{\mathcal{P}(x\mid y=0)}$ (7)
$=\left(\frac{\mathcal{P}(y=1\mid x)\mathcal{P}(x)}{\mathcal{P}(y=1)}\right)\left(\frac{\mathcal{P}(y=0\mid x)\mathcal{P}(x)}{\mathcal{P}(y=0)}\right)^{-1}$
$=\frac{\mathcal{P}(y=0)}{\mathcal{P}(y=1)}\frac{\mathcal{P}(y=1\mid x)}{\mathcal{P}(y=0\mid x)}.$
Equal prior probabilities are assumed for the source distributions, specifically $\mathcal{P}(y=0)=\mathcal{P}(y=1)$. This assumption is strategically made to prevent the estimation process from being skewed towards any particular class. By setting the prior probabilities to be equal, the ratio of marginal probabilities effectively cancels out, leaving:
@@
$\sigma^{-1}(z)=\ln\left(\dfrac{z}{1-z}\right),$ (9)
where $z$ represents any probability. The logit function has the property that the log odds of two probabilities equal the log of their ratio. In our case, applying the logit function to both the numerator and denominator of the expression for (8) results in the log ratio of the original probabilities:
$\ln\left(\dfrac{\mathcal{P}(y=1\mid x)}{\mathcal{P}(y=0\mid x)}\right)$ (10)
$=\sigma^{-1}\left(\mathcal{P}(y=1\mid x)\right)-\sigma^{-1}\left(\mathcal{P}(y=0\mid x)\right)$
@@ 3 Methodology
This research uses a neural network (NN) classifier, denoted $D_{\theta}$ with parameters $\theta$, to approximate the posterior probability $P(y=1\mid x)$. This network acts as a discriminator to classify input samples originating from the $p(x)$ or $q(x)$ distributions. The network architecture is designed to capture the discriminative features that distinguish these distributions effectively. In this study, we aim to understand how input data variations influence the ratio estimation between distributions. To isolate the impact of input data, we deliberately chose not to fine-tune the model across different experiments, maintaining a consistent architecture throughout. This approach allows us to examine the robustness of the estimator across various data types without the confounding effects of model optimization. The model parameters were selected to balance interpretability and provide meaningful insights, ensuring that our results illustrate the estimator’s practical capabilities rather than achieving optimal performance in each scenario. This strategy emphasizes the versatility and applicability of the estimator in real-world contexts, offering valuable insights into the challenges of synthetic data validation.
@@
$r_{\theta}(x)$ $=\exp\big{[}\sigma^{-1}\left(D_{\theta}(x)\right)\big{]}$ (13)
$\approx\exp\big{[}\sigma^{-1}(\mathcal{P}(y=1\mid x))\big{]}=r^{*}(x).$
@@
$\mathcal{L}(\theta)$ (14)
$=-\mathbb{E}_{p(x)}\big{[}\log\mathcal{D}_{\theta}(x)\big{]}-\mathbb{E}_{q(x)}\big{[}\log(1-\mathcal{D}_{\theta}(x))\big{]}.$
Once trained, the class probability estimator can construct MC estimates of the KL and JS divergences. The KL divergence between $p(x)$ and $q(x)$ can be estimated as
$D_{\mathbb{KL}}\big{(}p(x)\|q(x)\big{)}$ $=\mathbb{E}_{p(x)}\big{[}\log r^{*}(x)\big{]}$ (15)
$\approx\frac{1}{L}\sum_{i=1}^{L}\log r^{*}(x_{i})\approx\frac{1}{L}\sum_{i=1}^{L}\log r_{\theta}(x_{i})$
$=\frac{1}{L}\sum_{i=1}^{L}\sigma^{-1}\big{(}\mathcal{D}_{\theta}\left(x_{i}\right)\big{)},$
@@
$\approx$ $\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}2\mathcal{D}_{\theta}(x_{i})\big{)}$
$+\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}2-2\mathcal{D}_{\theta}(\tilde{x}_{i})\big{)}$
$=$ $\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}\mathcal{D}_{\theta}(x_{i})\big{)}$
$+\frac{1}{2L}\sum_{i=1}^{L}\log\big{(}1-\mathcal{D}_{\theta}(\tilde{x}_{i})\big{)},$
@@
$2\cdot D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}-\log 4$ (18)
$\geq\sup_{\theta}\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}+\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}.$
Following this analysis, we can arrive at the desired inequality. However, we can manipulate the inequality by negating both sides for a more convenient form.
$-2\cdot D_{\mathbb{JS}}\big{(}p(x)\|q(x)\big{)}+\log 4$ (19)
$\leq-\sup_{\theta}\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}+\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}$
$=\inf_{\theta}-\mathbb{E}_{p(x)}\big{[}\log D_{\theta}(x)\big{]}-\mathbb{E}_{q(x)}\big{[}\log(1-D_{\theta}(x))\big{]}$
$=\inf_{\theta}\mathcal{L}(\theta).$
@@
Figure 2: Generative model for divergence estimation between real and synthetic data. The generative model learns an approximation of $N$ samples from real data $x_{r}$, denoted $x_{s}$. Subsequently, $M$ samples are drawn from each distribution to train the divergence estimation discriminator. Finally, $L$ samples from each distribution estimate the divergence between real and synthetic data.
This study evaluates the dissimilarity between real and synthetic data generated by a Generative Model (GM). Minimizing the divergence between the real and synthetic data distributions is crucial in this context. Ideally, the divergence should approach zero, indicating that the generated data become virtually indistinguishable from the real data. This interchangeability allows the use of synthetic data in various applications where real data might be scarce or sensitive. From the discriminator’s perspective, achieving minimal divergence implies that it cannot reliably differentiate between real and synthetic samples, indicating successful data generation. For GMs, the distributions of interest become:
@@
Training data consist of $N$ samples drawn from the real data distribution, denoted $x_{r}\sim p_{R}(x)$. These data are used to train the GM, as illustrated in Fig. 2. The GM learns to approximate the real data distribution as $p_{S}(x)$, allowing us to sample from this distribution and obtain synthetic data points $x_{s}\sim p_{S}(x)$. Following the architecture described in Fig. 1, these synthetic samples are then used along with the real data samples to train the discriminator and estimate the divergence between the real and synthetic data distributions.